# Subspace Selected Variational Quantum Configuration Interaction with a Partial Walsh Series

### v1

@@ Abstract
Estimating the ground-state energy of a quantum system is one of the most promising applications for quantum algorithms. Here we propose a variational quantum eigensolver (VQE) Ansatz for finding ground state configuration interaction (CI) wavefunctions. We map CI for fermions to a quantum circuit using a subspace superposition, then apply diagonal Walsh operators to encode the wavefunction. The algorithm can be used to solve both full CI and selected CI wavefunctions, resuling in exact and near-exact solutions for electronic ground states. Both the subspace selection and wavefunction Ansatz can be applied to any Hamiltonian that can be written in a qubit basis. The algorithm bypasses costly classical matrix diagonalizations, which is advantageous for large-scale applications. We demonstrate results for several molecules using quantum simulators and hardware.
Introduction.– The energy eigenvalue estimation of many-body systems is a computational challenge inspiring continued improvements in both quantum and classical algorithms [1, 2, 40, 9, 36, 39, 48, 51, 41]. In strongly correlated regimes classical calculations can be limited by the growth of the parameter space or many-body entanglement [66, 38, 23, 50]. Quantum algorithms offer asymptotic advantages by natively representing quantum states, but may require large-scale error-correcting or fault-tolerant quantum computers to be competative platforms for routine eigenvalue estimation [33].
@@
In this Letter, we propose a diagonal Ansatz for VQE, which controls overparametrization of the state and avoids barren plateaus. Notably, the number of one- and two-qubit gates scales linearly with the number of SDs. There are two main components of the algorithm: superposition preparation and Ansatz application. First, on a quantum circuit we prepare a superposition over selected physically-relevant bitstrings. Second, we apply a diagonal Ansatz which effectively encodes the wavefunction coefficients in the appropriate bitstrings. The diagonal operator requires one ancilla qubit, and results in a probabalistic algorithm; however, the success probability is always 50%. Following standard VQE, we estimate the energy of the quantum circuit, and update the coefficients classically. We compute the ground-state energy for a range of molecules with IBMQ’s Torino processor and simulator. Our results provide a path forward for systematically improvable, and non-overparameterized Ansätze for many-body Hamiltonians, with relatively simple and physically motivated quantum circuits.
Theory and Methods.– A molecular electronic Hamiltonian under the Born-Oppenheimer approximation can be expressed as [65],
@@
Results.– We find the ground states of various molecules using VQE with either FCI or SCI, and benchmark the simulations with classical FCI and coupled-cluster single-doubles with perturbative triples (CCSD(T)) using PySCF [65, 13, 64]. We obtain the electronic Hamiltonian from OpenFermion, transformed to the qubit basis using the Jordan-Wigner transformation [73, 42], and use Qiskit to construct the quantum circuits for the statevector simulator, quantum device simulator, and IBMQ’s Torino quantum hardware, with details specified in the End Matter [64, 30, 3]. To estimate the energy we use the Hadamard test to measure the expectation values of Pauli strings, requiring another ancilla qubit [4]. For H2, we converge the energy to chemical accuracy, $1.6\times 10^{-3}$ Hartree, and H6 we converge the energy to $10^{-5}$ Hartree, and use the COBYLA and BFGS optimizers implemented in SciPy respectively [71]. For the device and noisy simulator we sample with 1024 measurements. Finally, we found that oversampling the WFT by $2D$ or $3D$ was sufficient to generate a full rank transformation, with the exact values of oversampling constants listed in the End Matter.
@@
For large-scale applications and molecules in non-minimal basis sets, it is advantageous to avoid the classical QR decomposition. As described above we randomly oversample the Walsh basis with $D\textrm{log}D$ functions. This tradeoff between the surjectivity of the Ansatz and classical resource efficiency provides a practical approach for systems which are not possible with a QR decomposition. As an example we use H2O in a 6-31G basis set resulting in 44 circuit parameters for 16 SDs, shown in Figure 3 [20]. We show results from the statevector simulator (teal squares) and noiseless Aer sampler (purple diamonds) using 10 qubits and $2^{17}$ shots.
@@
In general, our resource requirements are similar to those of recent high-quality ADAPT-VQE circuits [6, 54]. Here, we find the exact FCI solution under three different cases: spin and particle-number conserving, particle-number conserving, and no symmetry. The Walsh Ansatz has favorable scaling in both the two-qubit CNOT gates, and the number of variational parameters required for the one-qubit rotation gates. The number of CNOT gates for the Walsh Ansatz scales linearly with the number of determinants chosen, $\mathcal{O}(D)$. Furthermore, the Walsh Ansatz has the same number of variational parameters for $R_{Z}$ gates as the number of SDs in the selected expansion. This highlights that the Ansatz introduces no overparameterization of the chosen subspace, which helps in limiting spurious local minima in the optimization.
@@
Conclusions and Discussion.– Here, we present a variational quantum algorithm to estimate the ground-state energies of molecular Hamiltonians. This algorithm has two key steps, the first is preparing a state of physically-relevant bitstrings and the second is transforming the amplitudes of each SD with the truncated set of Walsh operators. After selecting the $D$ desired CI coefficients, we increase the parameter space to generate a full-rank transformation. A full-rank transformation can be generated by including $D\textrm{log}D$ randomly selected Walsh functions for the partial WFT [69]. The wavefunction Ansatz can be prepared with $\mathcal{O}(D\textrm{log}D)$ CNOT gates, and a $\mathcal{O}(D^{2}\textrm{log}^{2}D)$ classical operation. Furthermore, the Ansatz does not form a 2-design because it is diagonal and commutative, which mitigates the emergence of barren plateaus [41, 53, 52].
@@
Acknowledgements.– KHM acknowledges start-up funding from the University of Minnesota. ARS acknowledges a National Science Foundation (NSF) REU summer fellowship (CHE-2349246). All authors acknowledge the computational resources provided by the Minnesota Supercomputing Institute (MSI) at the University of Minnesota. The authors acknowledge the use of IBM Quantum Credits for this work.
Code and Data Availability.– The data from all results, along with the Python code to generate the data, are available at the Zenodo data repository [10].
Appendix A End Matter
Symmetry H6 H8 LiH BeH2 NH3 H2O
Spin & number 2 2 3 2 4 14
Number 2 2 3 2 4 14
Table 3: The oversampling constants to generate a full-rank partial Walsh-Fourier transform under spin and particle number and only particle number conservations.

### v2

@@ Abstract
Estimating the ground-state energy of a quantum system is one of the most promising applications for quantum algorithms. Here we propose a variational quantum eigensolver (VQE) Ansatz for finding ground state configuration interaction (CI) wavefunctions. We map CI for fermions to a quantum circuit using a subspace superposition, then apply diagonal Walsh operators to encode the wavefunction. The algorithm can be used to solve both full CI and selected CI wavefunctions, resulting in exact and near-exact solutions for electronic ground states. Both the subspace selection and wavefunction Ansatz can be applied to any Hamiltonian that can be written in a qubit basis. The algorithm bypasses costly classical matrix diagonalizations, which is advantageous for large-scale applications. We demonstrate results for several molecules using quantum simulators and hardware.
I Introduction
The energy eigenvalue estimation of many-body systems is a computational challenge inspiring continued improvements in both quantum and classical algorithms [1, 2, 40, 9, 36, 39, 48, 51, 41]. In strongly correlated regimes classical calculations can be limited by the growth of the parameter space or many-body entanglement [66, 38, 23, 50]. Quantum algorithms offer asymptotic advantages by natively representing quantum states, but may require large-scale error-correcting or fault-tolerant quantum computers to be competative platforms for routine eigenvalue estimation [33].
@@ I Introduction
In this Article, we propose a diagonal Ansatz for VQE, which controls overparametrization of the state and avoids barren plateaus. Notably, the number of one- and two-qubit gates scales linearly with the number of SDs. There are two main components of the algorithm: superposition preparation and Ansatz application. First, on a quantum circuit we prepare a superposition over selected physically-relevant bitstrings. Second, we apply a diagonal Ansatz which effectively encodes the wavefunction coefficients in the appropriate bitstrings. The diagonal operator requires one ancilla qubit, and results in a probabalistic algorithm; however, the success probability is always 50%. Following standard VQE, we estimate the energy of the quantum circuit, and update the coefficients classically. We compute the ground-state energy for a range of molecules with IBMQ’s Torino processor and simulator. Our results provide a path forward for systematically improvable, and non-overparameterized Ansätze for many-body Hamiltonians, with relatively simple and physically motivated quantum circuits.
II Theory and Methods
A molecular electronic Hamiltonian under the Born-Oppenheimer approximation can be expressed as [65],
@@ II Theory and Methods
III Results
We find the ground states of various molecules using VQE with either FCI or SCI, and benchmark the simulations with classical FCI and coupled-cluster single-doubles with perturbative triples (CCSD(T)) using PySCF [65, 13, 64]. We obtain the electronic Hamiltonian from OpenFermion, transformed to the qubit basis using the Jordan-Wigner transformation [73, 42], and use Qiskit to construct the quantum circuits for the statevector simulator, quantum device simulator, and IBMQ’s Torino quantum hardware, with details specified in the Appendix A [64, 30, 3]. To estimate the energy we use the Hadamard test to measure the expectation values of Pauli strings, requiring another ancilla qubit [4]. For H2, we converge the energy to chemical accuracy, $1.6\times 10^{-3}$ Hartree, and H6 we converge the energy to $10^{-5}$ Hartree, and use the COBYLA and BFGS optimizers implemented in SciPy respectively [71]. For the device and noisy simulator we sample with 1024 measurements. We find a full rank WFT by oversampling the Walsh basis and use the QR decomposition to find a $D$ by $D$ WFT.
@@ III Results
For large-scale applications and molecules in non-minimal basis sets, it is advantageous to avoid the classical QR decomposition. As described above we randomly oversample the Walsh basis with $D\textrm{log}D$ functions. This tradeoff between the surjectivity of the Ansatz and classical resource efficiency provides a practical approach for systems which are not possible with a QR decomposition. As an example we use H2O in a 6-31G basis set resulting in 44 circuit parameters for 16 SDs, shown in Figure 3 [20]. We show results from the statevector simulator (teal squares) and noiseless Aer sampler (purple diamonds) using 10 qubits and $2^{18}$ shots.
@@
We also investigate whether all $D\textrm{log}D$ oversampled Walsh functions are necessary for a full rank transformation. Figure 4 shows the likelihood of finding a full rank transformation by random selection as a function of the fraction of $D\textrm{log}D$ for several molecules. To generate the data, we perform 50 random selections of Walsh functions from the uniform distribution and find the rank of the transform, and compute the success probability as the computed rank divided by the full rank. We find that for all molecules we can find a full rank transformation by random selection at about 50% of $D\textrm{log}D$. As the system size increases, a smaller fraction of $D\textrm{log}D$ Walsh functions is required to achieve a full rank transformation, which agrees with the previously reported $\mathcal{O}(1-D^{-1})$ results [69]. As expected, as system size increases, the required oversampling rate becomes smaller.
Figure 4: The median success probability of obtaining a full-rank WFT for several molecules, using fractions of $D\textrm{log}D$ randomly selected Walsh functions. Error bars represent the standard deviation of 50 trials.
@@
In general, our resource requirements are similar to those of recent high-quality ADAPT-VQE circuits [6, 54]. Here, we find the exact FCI solution under three different cases: spin and particle-number conserving, particle-number conserving, and no symmetry.
The Walsh Ansatz has favorable scaling in both the two-qubit CNOT gates, and the number of variational parameters required for the one-qubit rotation gates with respect to the size of the parameter space. The number of CNOT gates for the Walsh Ansatz scales linearly with the number of determinants chosen, $\mathcal{O}(D)$. Furthermore, the Walsh Ansatz has the same number of variational parameters for $R_{Z}$ gates as the number of SDs in the selected expansion. This highlights that the Ansatz introduces no overparameterization of the chosen subspace, which helps in limiting spurious local minima in the optimization.
@@
IV Conclusions and Discussion
Here, we present a variational quantum algorithm to estimate the ground-state energies of molecular Hamiltonians. This algorithm has two key steps, the first is preparing a state of physically-relevant bitstrings and the second is transforming the amplitudes of each SD with the truncated set of Walsh operators. After selecting the $D$ desired CI coefficients, we increase the parameter space to generate a full-rank transformation. A full-rank transformation can be generated by including $D\textrm{log}D$ randomly selected Walsh functions for the partial WFT [69]. The wavefunction Ansatz can be prepared with $\mathcal{O}(D\textrm{log}D)$ CNOT gates, and a $\mathcal{O}(D^{2}\textrm{log}^{2}D)$ classical operation. Furthermore, the Ansatz does not form a 2-design because it is diagonal and commutative, which mitigates the emergence of barren plateaus [41, 53, 52].
@@ IV Conclusions and Discussion
V Acknowledgements
KHM acknowledges start-up funding from the University of Minnesota. ARS acknowledges a National Science Foundation (NSF) REU summer fellowship (CHE-2349246). All authors acknowledge the computational resources provided by the Minnesota Supercomputing Institute (MSI) at the University of Minnesota. The authors acknowledge the use of IBM Quantum Credits for this work.
VI Code and Data Availability
The data from all results, along with the Python code to generate the data, are available at the Zenodo data repository [10].
Appendix A Appendix: Device detail